# LaFAM: Unsupervised Feature Attribution with Label-free Activation Maps

### v1

@@ 4 Experiment
This evaluation is conducted using segmentation masks derived from the validation sets of ImageNet-S [14] and the PASCAL VOC 2012 [12, 11] datasets. To ensure a fair and consistent comparison, all methods employ nearest neighbor interpolation for upscaling the saliency maps. Furthermore, we isolate pertinent samples by selecting distinct categories and discarding masks with multiple classes. Specifically, we selected significant categories and discarded masks with multiple labels to reduce evaluative ambiguity. For the PASCAL VOC 2012 dataset, the image set was narrowed down to 13 labels. The refined datasets consisted of 11,294 and 753 samples from ImageNet-S and PASCAL VOC 2012, respectively.
@@
Table 1 presents the results for the ImageNet-1k validation set, comparing LaFAM with RELAX for SimCLR and SwAV in SSL settings, and with Grad-CAM in the supervised setting as a point of comparison. Similar to the ImageNet-S results, LaFAM demonstrates comparable performance to Grad-CAM and even slightly surpasses some of them and outperforms RELAX across all metrics as shown in Table 2.
In the SSL scenario, LaFAM outperforms RELAX across all metrics. Specifically, the higher Pointing-Game and Sparseness scores indicate more precise and focused saliency maps. RELAX shows a low Sparseness score, indicating that it tends to produce noticeably noisy saliency maps, especially for small objects. This observation is supported by the qualitative findings, as depicted in Figure 1 and Figure 4 of the Appendix.
In supervised settings, LaFAM is competitive with Grad-CAM across several metrics. As expected, Grad-CAM demonstrates superior results in the Sparseness metric. It implies that Grad-CAM generates less scattered explanations, as its design inherently highlights discriminative regions associated with specific class labels. However, CAM methods require labels and it can be a limiting factor since class scores are computed with respect to the most probable prediction. This can lead to wrong or misleading explanations when the prediction is incorrect. In contrast, a lower Sparseness score for LaFAM suggests that it captures a broader range of learned features, which can be beneficial compared to CAM methods, especially when an image contains multiple objects in a scene. It, therefore, may not be necessarily interpreted as a shortcoming but rather as a more universal approach as shown in Figure 2.
@@
Figure 3 shows examples of misclassifications on ImageNet-1k. It can be seen that LaFAM highlights the correct object compared to Grad-CAM. This indicates that the model has recognized the target object, but due to multiple objects in the scene, a wrong class may have a larger output. Hence, the saliency maps generated by Grad-CAM are focused on the wrong object.

### v2

@@ 4 Experiment
This evaluation is conducted using segmentation masks derived from the validation sets of ImageNet-S [14] and the PASCAL VOC 2012 [12, 11] datasets. To ensure a fair and consistent comparison, all methods employ nearest neighbor interpolation for upscaling the saliency maps.
Furthermore, we isolate pertinent samples by selecting distinct categories and discarding masks with multiple classes. For the PASCAL VOC 2012 dataset, the image set was narrowed down to 13 labels. The refined datasets consisted of 11,294 and 753 samples from ImageNet-S and PASCAL VOC 2012, respectively.
@@
Table 1 and Table 2 present the results for the ImageNet-1k and PASCAL VOC 2012 datasets, comparing LaFAM with RELAX for SimCLR and SwAV in SSL settings, and with Grad-CAM in the supervised setting as a point of comparison.
In the SSL scenario, LaFAM outperforms RELAX across all metrics for both datasets. Specifically, the higher Pointing-Game and Sparseness scores indicate more precise and focused saliency maps. RELAX shows a low Sparseness score, indicating that it tends to produce noticeably noisy saliency maps, especially for small objects. This observation is supported by the qualitative findings, as depicted in Figure 1 and Figure 4 in the Appendix.
In supervised settings, LaFAM is competitive with Grad-CAM across several metrics. As expected, Grad-CAM demonstrates superior results in the Sparseness metric. It implies that Grad-CAM generates less scattered explanations, as its design inherently highlights discriminative regions associated with specific class labels. However, CAM methods require labels and it can be a limiting factor since class scores are computed with respect to the most probable prediction. This can lead to wrong or misleading explanations when the prediction is incorrect.
In contrast, a lower Sparseness score for LaFAM suggests that it captures a broader range of learned features, which can be beneficial compared to CAM methods, especially when an image contains multiple objects in a scene. It, therefore, may not be necessarily interpreted as a shortcoming but rather as a more universal approach as shown in Figure 2.
@@
Figure 3 shows examples of misclassifications on ImageNet-1k. It can be seen that LaFAM accurately highlights the correct object, in contrast to Grad-CAM. This may indicate that the model has identified the target object, but the presence of multiple objects in the scene may lead to a higher output for an incorrect class. Consequently, the saliency maps produced by Grad-CAM are focused on the wrong object.